# Quantizing the Bosonic String on a Loop Quantum Gravity Background

### v1

@@ Abstract
With the goal of understanding whether or not it is possible to construct a string theory which is consistent with loop quantum gravity (LQG), we study alternate versions of the Nambu-Goto action for a bosonic string. We consider two types of modifications. The first is a phenomenological action based on the observation that LQG tells us that areas of two-surfaces are operators in quantum geometry and are bounded from below. This leads us to a string action which is similar to that of bimetric gravity. We provide formulations of the bimetric string for both the Nambu-Goto (second order) and Polyakov (first order) formulations. We explore the classical solutions of this action and its quantization and relate it to the conventional string solutions.
The second is an action in which the background geometry is described in terms of the pullback of the connection which describes the bulk geometry to the worldsheet. The resulting action is in the form of a gauged sigma model, where the spacetime co-ordinates are now vectors which transform under $ISO(D,1)$. We find that for the particular case of a constant background connection the action reduces to the bimetric action discussed above. We discuss classical solutions and quantization strategies for this action and its implications for the broader program of unifying string theory and loop quantum gravity.
@@ 1 Introduction
“Quantum gravity” refers to the broad enterprise dedicated to finding a complete, consistent theory in which quantum mechanics coexists peacefully with general relativity. There are two major approaches to quantum gravity which are widely recognized for their success in describing various aspects of physics at the Planck scale. We discuss each in turn below giving a bird’s eye view of research spanning a period of nearly sixty years.
@@
While this procedure is straightforward in principle, in practice it was impossible to implement in the quantum theory due to the highly complicated non-polynomial dependence of the diffeomorphism and Hamiltonian constraints on the configuration and momentum variables. This unfortunate state of affairs persisted until the late 1980s when Abhay Ashtekar [8, 9] recast general relativity as a theory of a densitized Minkowski tetrad
$\widetilde{E}^{\mu}_{I}=\frac{1}{\sqrt{-g}}e^{\mu}_{I}$ (1.3)
and the self-dual part of the $\mathfrak{sl}(2,\mathbb{C})$ spin-connection
$A_{\mu}^{IJ}=\omega^{IJ}_{\mu}-\frac{i}{2}\varepsilon^{IJ}_{\ \ KL}\omega^{KL}_{\mu}.$ (1.4)
Ashtekar realized that the constraints of general relativity, when expressed in terms of these “new variables”, simplified drastically and became polynomial functions of the tetrad and connection variables. It was quickly realized by various researchers that the resulting theory could be quantized using the same methods used to quantize Yang-Mills gauge theory [10, 11] and that the diffeomorphism constraint could be solved exactly in terms of so-called spin-network states [12, 13, 14].
@@
Before proceeding let us state our “central dogma”. We take as physically well-motivated the central result of LQG that at the Planck scale well-defined geometric operators - such as those that measure the area and volume of a given region of space - exist and have discrete spectra, thus the transition from no area/volume to some area/volume is a discrete jump rather than a continuous one. We now wish to understand how to construct a worldsheet action which reflects this central dogma.
The following is the outline of the paper. In section 2 we recall the Nambu-Goto action and suggest a simple modification which incorporates our central dogma, which is that there must exist a smallest possible area $\Delta$ which is physically measurable.  [^3] In recent work [27] a very similar expression to ours for the modified Nambu-Goto action has been arrived at by considering a version of LQG called “area metric gravity”. We show that this action can be expanded in terms of a small coupling constant $k=l_{pl}/l_{s}$ where the Planck length - the length at which quantum geometric effects become important - $l_{pl}$ is much smaller than the string scale $l_{s}$ and therefore $k\ll 1$. We point out that the lowest order correction to the NG action can be interpreted in the context of bimetric gravity. In section 3 we expand upon this observation and suggest a new action for the “bimetric string”, along the lines of bimetric gravity. We find both, the Nambu-Goto and the Polaykov forms of this action. Furthermore, we show that the phenomenological correction $\Delta$ introduced earlier is related to the pullback of the Kalb-Ramond field $B_{\mu\nu}$ into the worldsheet.
In section 7 we describe how to construct an action for the bosonic string which encodes the geometry of the bulk spacetime in the language of the connection formulation of general relativity. We discuss possible avenues for quantizing this action and further study its implications for the relationship between string theory and LQG.
@@ 7 Holography From Classical Strings
$D_{a}X^{\mu}(\tau,\sigma)=\partial_{a}X^{\mu}+k\mathcal{A}_{a}^{\mu}{}_{\nu}X^{\nu},$ (7.2)
@@ 7.3 The Einstein-Cartan String
$T^{A}{}_{B}$ $:=\begin{pmatrix}0&b^{1}&b^{2}&b^{3}&p^{0}/l\\ b^{1}&0&j^{3}&-j^{2}&p^{1}/l\\ b^{2}&-j^{3}&0&j^{1}&p^{2}/l\\ b^{3}&j^{2}&-j^{1}&0&p^{3}/l\\ -\epsilon p^{0}/l&-\epsilon p^{1}/l&-\epsilon p^{2}/l&-\epsilon p^{3}/l&0\end{pmatrix}$
@@
$\mathcal{D}_{a}X^{A}(\tau,\sigma)=\partial_{a}X^{A}+k\mathcal{A}_{a}^{A}{}_{B}X^{B}.$ (7.9)
@@ 8 Discussion
We also briefly explored an alternative approach to incorporate LQG quantum geometric effects into string theory by considering Afshordi’s “emergent dimensions” model [28] with some key insights from LQG complementing said model, creating a more concrete picture of the mechanism behind the emerging dimensions.
@@ 8.2 Closed String Spin-1 States
Another candidate for what the vector excitations could relate to is the Ashtekar electric fields $\widetilde{E}^{\mu}_{I}$,

### v2

# A Loop Quantum Gravity Inspired Action for the Bosonic String and Emergent Dimensions at Large Scales

@@ Abstract
We propose a modification of the Nambu-Goto action for the bosonic string which is compatible with the existence of a minimum area at the Planck scale. The result is a phenomenological action based on the observation that LQG tells us that areas of two-surfaces are operators in quantum geometry and are bounded from below. This leads us to a string action which is similar to that of bimetric gravity. We provide formulations of the bimetric string action for both the Nambu-Goto (second order) and Polyakov (first order) formulations. We explore the classical solutions of this action and its quantization and relate it to the conventional string solutions.
We further construct a string action in which the effect of the background geometry is described in terms of the pullback of the bulk connection, which encodes the bulk geometry, to the worldsheet. The resulting string action is in the form of a gauged sigma model, where the spacetime co-ordinates are now vectors which transform under the Poincaré group $ISO(D,1)$. This requires the introduction of an auxiliary bulk co-ordinate which has a natural interpretation as a holographic or scale direction. We discuss possible cosmological implications of such a large scale emergent dimension.
@@ 1 Introduction
“Quantum gravity” refers to the broad enterprise dedicated to finding a complete, consistent theory in which quantum mechanics coexists peacefully with general relativity. There are two major approaches to quantum gravity which are widely recognized for their success in describing various aspects of physics at the Planck scale.
@@
In this paper we present a phenomenological modification of the Nambu-Goto action for the bosonic string which has the property that the worldsheet area can never be less than some fixed constant value $\Delta$ which represents the minimum allowed quantum of area predicted by LQG.
The following is the outline of the paper along with a brief summary of what is accomplished in each section. For those seeking a rapid review of the principle areas of diffence between the two theories subsection 1.1 provides a bird’s eye view of research spanning a period of nearly sixty years.
Those already familiar with the essentials of LQG (primarily area quantization) and the question of background independence of string theory can skip ahead to section 2 where we being by recalling the Nambu-Goto action and suggest a simple modification which incorporates our central dogma, which is that there must exist a smallest possible area $\Delta$ which is physically measurable.  [^1] In recent work [8] a very similar expression to ours for the modified Nambu-Goto action has been arrived at by considering a version of LQG called “area metric gravity”. We show that this action can be expanded in terms of a small coupling constant $k=l_{pl}/l_{s}$ where the Planck length - the length at which quantum geometric effects become important - $l_{pl}$ is much smaller than the string scale $l_{s}$ and therefore $k\ll 1$. We point out that the lowest order correction to the NG action can be interpreted in the context of bimetric gravity. In section 3 we expand upon this observation and suggest a new action for the “bimetric string”, along the lines of bimetric gravity. We find both, the Nambu-Goto and the Polaykov forms of this action. Furthermore, we show that the phenomenological correction $\Delta$ introduced earlier is related to the pullback of the Kalb-Ramond field $B_{\mu\nu}$ into the worldsheet.
In section 7 we describe how to construct an action for the bosonic string which encodes the geometry of the bulk spacetime in the language of the connection formulation of general relativity. We find that this can be done by viewing the worldsheet embedding co-ordinates $X$’s as living in the fundamental representation of a suitable gauge group $G$. Consistency between the dual roles played by the $X$’s as both co-ordinates for the background manifold and and as objects which can be parallel transported using a connection, requires that the gauge group $G$ be the Poincaré group $ISO(D+1,1)$. In other words the string which originally was embedded in a $D+1$ dimensional background, now lives in a $D+2$ dimensional geometry which has one extra “emergent” dimension which has the natural interpretation as a scaling dimension. We therefore view our construction as providing a natural framework for classical holography, where the scale dimension is already built into the theory at the classical level instead of something which arises due to non-perturbative physics.
1.1 Historical Background
@@
While this procedure is straightforward in principle, in practice it was impossible to implement in the quantum theory due to the highly complicated non-polynomial dependence of the diffeomorphism and Hamiltonian constraints on the configuration and momentum variables. This unfortunate state of affairs persisted until the late 1980s when Abhay Ashtekar [10, 11] recast general relativity as a BF theory of a dualised $\mathfrak{sl}(2,\mathbb{C})$ connection (which becomes the self-dual part of said connection in the case of $\beta=i$)
$A_{\mu}^{IJ}=\omega^{IJ}_{\mu}-\frac{\beta}{2}\varepsilon^{IJ}_{\ \ KL}\omega^{KL}_{\mu}$ (1.3)
and its conjugate momenta
$B^{IJ}_{\mu\nu}=\left(\delta^{I}_{K}\delta^{J}_{L}+\frac{1}{2\beta}\varepsilon^{IJ}_{\ \ KL}\right)e^{K}_{[\mu}e^{L}_{\nu]},$ (1.4)
which when restricted to a 3-dimensional hypersurface where $*(e\wedge e)=0$ reduces to
$A^{i}_{a}=\omega^{i}_{a}+\beta K^{i}_{a}$ (1.5)
$E^{a}_{i}=\sqrt{q}e^{a}_{i}=\frac{1}{2}\varepsilon^{abc}\varepsilon_{ijk}e^{j}_{b}e^{k}_{c},$ (1.6)
with $K^{i}_{a}$ being the triad version of the extrinsic curvature $K_{ab}=\frac{1}{2}\mathcal{L}_{n}q_{ab}$ and $q$ the 3-metric. Ashtekar realized that the constraints of general relativity, when expressed in terms of these “new variables”, simplified drastically and became polynomial functions of the tetrad and connection variables. It was quickly realized by various researchers that the resulting theory could be quantized using the same methods used to quantize Yang-Mills gauge theory [12, 13] and that the diffeomorphism constraint could be solved exactly in terms of so-called spin-network states [14, 15, 16].
@@ 2 Modified Nambu Goto Action
Before proceeding let us restate our “central dogma”. We take as physically well-motivated the central result of LQG that at the Planck scale well-defined geometric operators - such as those that measure the area and volume of a given region of space - exist and have discrete spectra, thus the transition from no area/volume to some area/volume is a discrete jump rather than a continuous one. We now wish to understand how to construct a worldsheet action which reflects this central dogma.
@@ 7 Holography From Classical Strings
$D_{a}X^{\mu}(\tau,\sigma)=\partial_{a}X^{\mu}+\mu\mathcal{A}_{a}^{\mu}{}_{\nu}X^{\nu},$ (7.2)
@@ 7.3 The Einstein-Cartan String
$T^{A}{}_{B}$ $:=\begin{pmatrix}0&b^{1}&b^{2}&b^{3}&p^{0}/l\\ b^{1}&0&j^{3}&-j^{2}&p^{1}/l\\ b^{2}&-j^{3}&0&j^{1}&p^{2}/l\\ b^{3}&j^{2}&-j^{1}&0&p^{3}/l\\ \epsilon p^{0}/l&-\epsilon p^{1}/l&-\epsilon p^{2}/l&-\epsilon p^{3}/l&0\end{pmatrix}$
@@
$\mathcal{D}_{a}X^{A}(\tau,\sigma)=\partial_{a}X^{A}+\mu\mathcal{A}_{a}^{A}{}_{B}X^{B}.$ (7.9)
@@
7.4 Gauged Polyakov Equations of Motion
By varying the action (7.8), we get two sets of constraints from the variation of $g_{ab}$ and $A_{a}^{AB}$, namely the vanishing of the worldsheet energy-momentum and spin-linear-momentum tensors $T_{ab}$ and $\sigma_{a}^{AB}$
$T_{ab}:=\mathcal{D}_{a}X^{A}\mathcal{D}_{b}X^{B}G_{AB}-\\ -g_{ab}g^{cd}\mathcal{D}_{c}X^{A}\mathcal{D}_{d}X^{B}G_{AB}=0$ (7.13)
$\sigma^{AB}_{a}:=\mathcal{D}_{a}X^{[A}X^{B]}=0.$ (7.14)
The former allows us to solve for the worldsheet “covariant metric” $g_{ab}$ as
$g_{ab}=2f\mathcal{D}_{a}X^{A}\mathcal{D}_{b}X^{B}G_{AB}$ (7.15)
where
$\frac{1}{f}=g^{ab}\mathcal{D}_{a}X^{A}\mathcal{D}_{b}X^{B}G_{AB},$ (7.16)
while the latter should allow one to solve for $\mathcal{A}_{a}$ in terms of $X^{A}$.
The variation of the action w.r.t. $X^{A}$ results in the equation of motion of the embbeding fields
$\mathcal{D}_{a}[\sqrt{-g}g^{ab}\mathcal{D}_{b}X^{B}G_{BC}]=\\ =\sqrt{g}g^{ab}\mathcal{D}_{a}X^{A}\mathcal{D}_{b}X^{B}\partial_{C}G_{AB},$ (7.17)
which in worldsheet “covariant conformal gauge” $g^{ab}=\mathrm{e}^{-\phi}\eta^{ab}$ and for flat background geometry $G_{AB}=\eta^{\prime}_{AB}$ with
$[\eta^{\prime}_{AB}]=\begin{bmatrix}[\eta_{\mu\nu}]&0\\ 0&\epsilon\end{bmatrix}$ (7.18)
simplifies to
$\eta^{ab}\mathcal{D}_{a}\mathcal{D}_{b}X^{A}\eta^{\prime}_{AB}=0.$ (7.19)
From here on, we’ll work on the case of $\epsilon=-1$ (adS background), for which the above equation can be expanded as
$\eta^{ab}\big{(}\partial_{a}\partial_{b}X^{A}+2\mu A^{A}_{aB}\partial_{b}X^{B}+\\ +\mu\partial_{a}A^{A}_{bB}X^{B}+\mu^{2}A^{A}_{aB}A^{B}_{bC}X^{C}\big{)}=0.$ (7.20)
To produce the desired bulk geometry, the background connection can be taken as
$\mathcal{A}_{\mu}=\begin{bmatrix}0&0&0&0&\delta^{0}_{\mu}\\ 0&0&0&0&\delta^{1}_{\mu}\\ 0&0&0&0&\delta^{2}_{\mu}\\ 0&0&0&0&\delta^{3}_{\mu}\\ -\delta^{0}_{\mu}&\delta^{1}_{\mu}&\delta^{2}_{\mu}&\delta^{3}_{\mu}&0\end{bmatrix},$ (7.21)
from which the pulledback connection $\mathcal{A}_{a}=\partial_{a}X^{\mu}\mathcal{A}_{\mu}$ is thus
$\mathcal{A}_{a}=\begin{bmatrix}0&0&0&0&\partial_{a}X^{0}\\ 0&0&0&0&\partial_{a}X^{1}\\ 0&0&0&0&\partial_{a}X^{2}\\ 0&0&0&0&\partial_{a}X^{3}\\ -\partial_{a}X^{0}&\partial_{a}X^{1}&\partial_{a}X^{2}&\partial_{a}X^{3}&0\end{bmatrix},$ (7.22)
whereby imposing conformal gauge on the induced metric $\gamma_{ab}=\partial_{a}X^{\mu}\partial_{b}X^{\nu}\eta_{\mu\nu}=\mathrm{e}^{\varphi}\eta_{ab}$ we can separate equations for the Minkowski sector $X^{\mu}$ and for the holographic dimension $X^{r}$:
$\eta^{ab}\partial_{a}\partial_{b}X^{\mu}+\\ +\frac{\mu\eta^{ab}}{1+\mu X^{r}}\left(2\partial_{a}X^{r}+k\partial_{a}X^{\nu}X_{\nu}\right)\partial_{b}X^{\mu}=0$ (7.23)
$\eta^{ab}\partial_{a}\partial_{b}X^{r}+2\mu^{2}\mathrm{e}^{\varphi}X^{r}+4\mu\mathrm{e}^{\varphi}-\\ -\frac{\mu^{2}\eta^{ab}}{1+\mu X^{r}}\left(2\partial_{a}X^{r}+\mu\partial_{a}X^{\nu}X_{\nu}\right)\partial_{b}X^{\mu}X_{\mu}=0.$ (7.24)
Assuming the perturbations on the string are small $|X^{A}|\ll 1$, we get a massive simplification
$\eta^{ab}\partial_{a}\partial_{b}X^{\mu}=0$ (7.25)
$\eta^{ab}\partial_{a}\partial_{b}X^{r}+2\mu^{2}\mathrm{e}^{\varphi}X^{r}=-4\mu\mathrm{e}^{\varphi}.$ (7.26)
Thus, the Minkowski sector behaves like regular ST for small perturbations, and for the holographic dimension we can solve using a Green’s function
$X^{r}(x)=-4\mu\int\mathrm{d}^{2}x^{\prime}G(x;x^{\prime})\mathrm{e}^{\varphi(x^{\prime})},$ (7.27)
with the Green’s function satisfying
$(\partial_{\sigma}^{2}-\partial_{\tau}^{2}+2\mu^{2}\mathrm{e}^{\varphi})G(x;x^{\prime})=\delta^{2}(x,x^{\prime}).$ (7.28)
To find this Green’s function, we must fix the conformal factor to be of the form $\mathrm{e}^{\varphi}=f(\tau)-g(\sigma)$, thus allowing us to separate variables in the homogeneous case of eq (7.28) into
$X^{\prime\prime}-2\mu^{2}g(\sigma)X=-\lambda X$ (7.29)
$\ddot{T}-2\mu^{2}f(\tau)T=-\lambda T.$ (7.30)
These are Schrodinger’s equations for potentials given by $\mu f(\tau)$ and $\mu g(\sigma)$, as such if we assume that $f$ and $g$ are harmonic oscillator potentials $f(\tau)=\omega^{2}\tau^{2}/2$ and $g(\sigma)=k^{2}\sigma^{2}/2$ we can find known solutions for separation constant $\lambda=2\mu E=\mu k(2n+1)=\mu\omega(2n+1)$, so by the Spectral Theorem our separated spatial and temporal Green’s functions are
$X(\sigma,\sigma^{\prime})=\sum_{n\geq 0}-\frac{X_{n}(\sigma)X_{n}(\sigma^{\prime})}{\mu k(2n+1)}$ (7.31)
$T(\tau,\tau^{\prime})=\sum_{m\geq 0}-\frac{T_{m}(\tau)T_{m}(\tau^{\prime})}{\mu\omega(2m+1)},$ (7.32)
where $X_{n}$ and $T_{m}$ are harmonic oscillator eigenfunctions. With this, the full Green’s function $G=T.X$ and the solution for $X^{r}$ are thus
$G(x;x^{\prime})=\sum_{m\geq 0}\sum_{n\geq 0}\frac{T_{m}(\tau)T_{m}(\tau^{\prime})}{\mu\omega(2m+1)}\frac{X_{n}(\sigma)X_{n}(\sigma^{\prime})}{\mu k(2n+1)}$ (7.33)
$X^{r}(x)=-4\mu\sum_{m\geq 0}\sum_{n\geq 0}\frac{T_{m}(\tau)}{\mu\omega(2m+1)}\frac{X_{n}(\sigma)}{\mu k(2n+1)}\times\\ \times\int\mathrm{d}^{2}x^{\prime}T_{m}(\tau^{\prime})X_{n}(\sigma^{\prime})\left(\frac{\omega^{2}\tau^{\prime 2}}{2}-\frac{k^{2}\sigma^{\prime 2}}{2}\right).$ (7.34)
@@ 8 Discussion
We also briefly explored an alternative approach to incorporate LQG quantum geometric effects into String Theory by considering Afshordi’s model [9] with some key insights from LQG complementing said model, creating a more concrete picture of the mechanism behind the emerging dimensions and how ultimately spacetime (and thus the true quantum gravitational field) might be actually $(2+1)$-dimensional at the Planck scale.
@@ 8.2 Closed String Spin-1 States
Another candidate for what the vector excitations could relate to is the Ashtekar electric field $\widetilde{E}^{\mu}_{I}$,